OpenStudy (anonymous):

solve logaritmo equation 1/2 log3 x= 2 log3 2

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle \frac{1}{2} \log_3 x= 2\log_3 2}\) like this ?

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle \frac{1}{2} \log_3 x= 2\log_3 2}\) \(\large\color{black}{\displaystyle \log_3 x= 4\log_3 2}\) \(\large\color{black}{\displaystyle \log_3 x= \log_3 \sqrt[4]{2}}\)

OpenStudy (solomonzelman):

then \(\large\color{black}{\displaystyle x=?}\)

OpenStudy (anonymous):

my book said 16

OpenStudy (solomonzelman):

oh, my bad, it should say, in my last line: \(\normalsize\color{royalblue}{ \rm \log_3x =\log_32^4 }\)